### v2

# Neural network for excess noise estimation in continuous-variable quantum key distribution under composable finite-size security

Abstract
Parameter estimation is a critical step in continuous-variable quantum key distribution (CV-QKD), as the statistical uncertainty from a finite data size leads to pessimistic worst-case bounds that drastically reduce the secret key rate and range. While machine learning techniques have been proposed for this task, they have lacked the rigorous statistical framework necessary for integration into a composable security proof. In this work, we bridge this gap by introducing a statistically rigorous framework for using neural networks for parameter estimation in CV-QKD with quantifiable composable security. We develop a neural network estimator for the excess noise and, crucially, derive its worst-case confidence interval using a delta method approach, ensuring the estimation fails with a probability not exceeding $\epsilon_{\rm PE}$. This allows the network to be integrated into a parameter estimation protocol that is operationally equivalent to the standard maximum likelihood method but yields significantly tighter parameter bounds. Our numerical results demonstrate that this method provides substantially more precise estimates, which directly translates into a higher secret key rate and extended transmission distance over a fiber channel under a collective Gaussian attack. This work establishes that machine learning can be securely and effectively harnessed to overcome a key performance limitation in practical CV-QKD systems.
quantum cryptography; continuous-variable quantum key distribution; parameter estimation; neural networks.
I Introduction
Security in communication is a fundamental aspect of contemporary society, as it enables the sharing of sensitive information without the risk of potential leaks Shamir (1979); United Nations (2024); European Union Agency for Cybersecurity (2024); Pirandola et al. (2020). However, the emergence of recent algorithms has threatened this security, such as the quantum algorithm for factoring integers in logarithmic time Shor (1994, 1999). Considering these challenges, efforts in the field of quantum communication have been made to use properties inherent to quantum physics to ensure unconditionally secure communication Gisin et al. (2002); Renner (2006); Portmann and Renner (2022); Scarani et al. (2009a). In this context, the use of continuous-variable quantum key distribution (CV-QKD) emerges as a potential alternative, since it has greater adaptability to the current components found in coherent optical telecommunications systems Usenko et al. (2025); Zhang et al. (2024); Laudenbach et al. (2018); Diamanti and Leverrier (2015).
In a generic CV-QKD protocol, quantum information can be encoded onto coherent states by modulating the amplitude and phase quadratures of laser light, typically using electro-optical modulators at the transmitter to establish a secret key between two legitimate QKD users (Alice and Bob) Grosshans and Grangier (2002a); Grosshans et al. (2003a); Weedbrook et al. (2004). These states are transmitted through a quantum channel that is assumed to be fully under the control of a potential eavesdropper, conventionally referred to as Eve. The security of CV-QKD protocols employing Gaussian-modulated coherent states was initially proven in the asymptotic limit Van Assche et al. (2005); Navascués et al. (2006); García-Patrón and Cerf (2006) and later extended to the finite-size regime, ensuring universal composability against both collective Leverrier et al. (2010) and general coherent attacks Furrer et al. (2012). Beyond continuous modulation, discrete modulation of coherent states has also been extensively studied and shown to offer promising performance and security guarantees Leverrier and Grangier (2009); Denys et al. (2021); Bäuml et al. (2024).
Since the channel is under the control of Eve, the smooth min-entropy has to be bounded by the worst case compatible with the observed measurement data Devetak and Winter (2005). Thus, a fundamental procedure in CV-QKD is the estimation of the channel parameters, such as the transmittance $T$ and the excess noise $\xi$ Usenko et al. (2025); Zhang et al. (2024); Laudenbach et al. (2018); Diamanti and Leverrier (2015). In principle, there are other parameters to be estimated, but transmittance and excess noise have the most significant impact on the secret-key rate, where the latter has a drastic impact for long distances Huang et al. (2016); Laudenbach et al. (2018); Hajomer et al. (2024). For a finite-key security analysis, parameter estimation must ensure that the key is secure against any eavesdropper attack, up to a probability of failure Scarani and Renner (2008); Cai and Scarani (2009).
There is broad consensus in the literature that the maximum likelihood estimation (MLE) method offers strong security guarantees for the protocol Thearle et al. (2016); Pirandola and Papanastasiou (2024); Papanastasiou et al. (2017); Zhang et al. (2020), as its confidence intervals can be explicitly computed and depend directly on the chosen significance level $\epsilon_{PE}$ Monfort (1982). In fact, the first finite-key analysis under the assumption of collective Gaussian attacks was provided in ref. Leverrier et al. (2010) using MLE. Recently, machine learning techniques have been introduced for various tasks in CV-QKD Long et al. (2023), including parameter estimation Liu et al. (2018); Chin et al. (2021); Luo et al. (2022). However, most of these works do not account for the probability of estimation failure when neural networks are employed. Consequently, there is currently no established statistical security framework for using neural networks within the CV-QKD context.
In this work, we provide a finite-size security analysis demonstrating that neural networks can be reliably used for CV-QKD with quantifiable failure probabilities $\epsilon_{PE}$, which possesses an operational interpretation and composability security. We demonstrated our analysis in a neural network architecture using a parameter estimation protocol (PEP) operationally equivalent to the standard method presented in ref. Leverrier et al. (2010). Our results showed that neural networks can provide more precise estimations in order to gain more distances without compromising the security of the protocol.
II Protocol and model definition
In this work, we will consider the coherent-state protocol with Gaussian modulation Grosshans and Grangier (2002b). The protocol starts with Alice preparing $N$ displaced vacuum states $\lvert q_{i}+ip_{i}\rangle$ by modulating both the amplitude and phase quadratures. The displacements $q_{i}$ and $p_{i}$ are independent random variables drawn from the normal distribution $\mathcal{N}(0,V_{A})$. These states are transmitted over an untrusted channel with transmission $T$ and excess noise $\xi$, assumed to be under Eve’s control. Upon reception, Bob performs homodyne detection by switching randomly between the phase and amplitude quadratures. Here, Alice’s use of modulation to prepare these states constitutes a prepare-and-measure protocol, known to have an entanglement-based equivalent Grosshans et al. (2003b), which is used for the security analysis García-Patrón and Cerf (2006).
For Gaussian channels, the relationship between the signal sent by Alice and the signal received by Bob is given by the linear model
$y_{i}=tx_{i}+z_{i}\,$, (1)
where $\{y_{i}\}_{N}$ and $\{x_{i}\}_{N}$ are the classical data related to the random variables of Bob and Alice, respectively. In the measurement, Alice’s signal is affected by the parameter $t=\sqrt{T}$ and the noise $\{z_{i}\}_{N}$, represented by a random variable with zero mean and variance $\sigma^{2}=\mu+t^{2}\xi$. The parameter $\mu$ is the quantum duty (“qu-duty”) associated with detection: $\mu=1$ for homodyne and $\mu=2$ for heterodyne Pirandola (2021). Operationally, the transmission can be evaluated with
$T\approx\eta_{\text{eff}}10^{-0.02d}$ (2)
where $\eta_{\text{eff}}$ is the known quantum efficiency of Bob’s detection and $d$ is the distance in kilometers between Alice and Bob. In this case, we assume an optical fiber with losses of $0.2$ dB per kilometer. More generally, we consider the possibility that Eve can access side-channel information resulting from imperfections in the detection setup. In this analysis, we focus on the receiver’s detection inefficiency and assume that the fraction $1-\eta_{\text{eff}}$ of the incoming photons that are not detected is not simply lost to the environment, but is instead collected by Eve and incorporated into her attack strategy Pirandola (2021).
In the context of QKD, the objective is to extract a positive secret key rate while guaranteeing composable security considering the signals $\{y_{i}\}_{N}$ and $\{x_{i}\}_{N}$ in the presence of channel loss and excess noise Scarani et al. (2009b); Renner (2008); Lo et al. (2014). A widely adopted method for evaluating this quantity is the Devetak–Winter bound Devetak and Winter (2005)
$I(x:y)-\sup_{\mathcal{N}:A^{\prime}\rightarrow B}\chi(y:E)$ (3)
where $I(x:y)$ denotes the mutual information between Alice’s and Bob’s classical variables $x$ and $y$ Cover (1999), while $\chi(y:E)$ represents the Holevo information between Bob’s variable $y$ and the adversary’s quantum system $E$ Nielsen and Chuang (2010). The supremum is taken over all channels $\mathcal{N}:A\rightarrow B$ that are consistent with the statistics observed by Alice and Bob during the parameter estimation. For the entanglement-based protocol Grosshans et al. (2003b), it can be described by the covariance matrix
$\Gamma=\begin{pmatrix}(V_{A}+1)\mathbb{I}_{2}&tZ\sigma_{z}\ \\ tZ\sigma_{z}&(t^{2}V_{A}+\sigma^{2})\mathbb{I}_{2}\\ \end{pmatrix}\,$, (4)
where $\sigma_{z}$ is the Pauli matrix and $Z=\sqrt{V_{A}^{2}+2V_{A}}$ for Gaussian modulation Laudenbach et al. (2018).
Using the covariance matrix, $\chi(y:E)$ can be determined by its symplectic eigenvalues Laudenbach et al. (2018). Following the quantum stage, classical data processing and a mathematically rigorous security analysis are performed to distill a secret key of certified length Pirandola (2021). However, Eq (3) does not take into account the effects of post-processing data after Bob’s measurements. For example, the parameter estimation significantly impacts this value because $m$ signals are used for this estimation, reducing the total number of data used to generate the raw key to $n\equiv N-m$ Cai and Scarani (2009). Also, the reconciliation efficiency $\beta$ is another important value to be considered, since it estimates the amount of information Bob can recover from Alice, limiting the mutual information Almeida et al. (2023).
Furthermore, since the finite number of quantum states exchanged by Alice and Bob inevitably reduces the achievable key length, incorporating finite-size effects is indispensable to guarantee composable security Canetti (2001) up to a failure probability $\epsilon_{PE}$, thereby ensuring the protocol remains operationally meaningful in realistic conditions Jain et al. (2022). This can be quantified considering the statistical error present in post-processing steps Leverrier et al. (2010); Pirandola (2021), characterized by
$\epsilon=p_{\text{ec}}\epsilon_{\mathrm{PE}}+\epsilon_{\text{cor}}+\epsilon_{\text{sec}}\,$ (5)
where $\epsilon_{\text{cor}}$ and $\epsilon_{\text{sec}}$ indicate that the protocol satisfies $\epsilon$-correctness and $\epsilon$-secrecy, respectively. The secrecy parameter $\epsilon_{\text{sec}}$ can be further decomposed as $\epsilon_{\text{sec}}=\bar{\epsilon}+\epsilon_{\mathrm{PA}}$, where $\bar{\epsilon}$ is the smoothing parameter and $\epsilon_{\mathrm{PA}}$ denotes the probability of failure of the privacy amplification step. Additionally, $p_{\text{ec}}\epsilon_{PE}$ accounts for the probability of failure in parameter estimation, while $p_{\text{ec}}=1-\text{FER}$ represents the probability of successful error correction, with FER being the frame error rate Furrer et al. (2012); Pirandola (2021).
The parameter $\epsilon$ must be composable and have an operational interpretation in order to ensure that it meets the security requirements Renner (2006). In the context of DV-QKD, ref.  Scarani and Renner (2008) showed that $\epsilon$ satisfies the composability criterion and corresponds to the maximum failure probability of the protocol, meaning the maximum probability that an eavesdropper obtains non-negligible information about the final key $k_{\epsilon}$. This notion was later extended to CV-QKD in ref.  Leverrier et al. (2010). Specifically, the parameter $\epsilon_{\mathrm{PE}}$ affects the estimation of the covariance matrix in Eq. (4), requiring the replacement of the Holevo information $\chi(y:E)$ by its smooth version $\chi_{\epsilon_{\mathrm{PE}}}(y:E)$. The parameters $\epsilon_{\mathrm{PA}}$ and $\bar{\epsilon}$ enter the secret key rate expression via the finite-size correction term $\Delta(n)$, which adjusts the asymptotic key rate to account for statistical fluctuations and composable security requirements Renner (2006). Thus, the secret key rate is finally written as
$k_{\epsilon}=\frac{np_{EC}}{N}(\beta I(x:y)-\chi_{\epsilon_{PE}}(y:E)-\Delta(n))$ (6)
and the last term is explicitly defined as
$\Delta(n)\equiv 4\log_{2}(\sqrt{d}+2)\sqrt{\frac{1}{n}\log_{2}\left(\frac{18}{p_{\mathrm{ec}}^{2}\epsilon_{s}^{4}}\right)}+\frac{2}{n}\log_{2}(1/\epsilon_{PA})\,$, (7)
where $d$ denotes the number of bits per quadrature used during discretization, $\bar{\epsilon}$ is the smoothing parameter, and $\epsilon_{\mathrm{PA}}$ is the failure probability of the privacy‐amplification procedure Leverrier et al. (2010). Both $\bar{\epsilon}$ and $\epsilon_{\mathrm{PA}}$ are intermediate quantities to be optimized numerically. The first term of $\Delta(n)$, namely the square‐root term, quantifies the convergence rate of the smooth min‐entropy — the relevant metric for key length — of an i.i.d. state under collective attacks toward its von Neumann entropy, since only in the asymptotic limit does the smooth min‐entropy equal the von Neumann entropy. Its derivation was done in ref. Pirandola (2021), which used the framework proposed in ref. Tomamichel (2013). The second term directly reflects the security contribution of the failure probability $\epsilon_{\mathrm{PA}}$ in the privacy‐amplification step.
III Parameter estimation effects on protocol security
The parameter estimation is, without any doubt, the main problem for CV-QKD in finite-size scenario: The uncertainty related to the estimation limits the secret key-rate, since one can never estimate a secret key-rate below its real value Leverrier et al. (2010). From the covariance matrix in Eq. (4), one can see that we need to compute both the transmittance and excess noise values. In this post-processing pipeline, we consider that Alice computes the covariance matrix with the data Bob publishes on the authenticated channel, so the variables $V_{A}$ and $Z$ are not considered problematic in the parameter estimation stage.
The main problem here is that Bob and Alice do not know these parameters, since it is assumed that the channel can be freely controlled by Eve. The law of large numbers guarantees that when $m\rightarrow\infty$ $\mathbb{E}[\hat{t}]\equiv t$ and $\mathbb{E}[\hat{\sigma}^{2}]\equiv\sigma^{2}$ Casella and Berger (2024), so there is no need for error analysis in asymptotic limit. In practical implementations, it is obviously impossible to achieve this result, such that one needs to consider the probability of failure of the parameter estimation $\epsilon_{PE}$ using statistical analysis for the estimators.
The maximum likelihood estimation (MLE) is widely recognized as the standard method in the field, since it is compatible with statistical analyses considering the confidence interval Thearle et al. (2016); Pirandola and Papanastasiou (2024); Papanastasiou et al. (2017); Zhang et al. (2020). For the linear model,
$\hat{t}=\sum_{i}^{m}\frac{y_{i}x_{i}}{x_{i}^{2}}\,\quad\text{and}\,\quad\hat{\sigma}^{2}=\sum_{i}^{m}\frac{(y_{i}-\hat{t}x_{i})^{2}}{m}\,.$ (8)
The confidence interval is then computed considering the lower bound for $t_{min}$ and the upper bound for $\sigma^{2}_{max}$:
$t_{min-\text{MLE}}\approx\hat{t}_{\text{MLE}}-z_{\epsilon_{PE}/2}\sqrt{\frac{\hat{\sigma}^{2}}{mV_{A}}}\,$, (9)
$\sigma^{2}_{\max-\text{MLE}}\approx\hat{\sigma}^{2}_{MLE}+z_{\epsilon_{PE}/2}\frac{\hat{\sigma}^{2}\sqrt{2}}{\sqrt{m}}\,.$ (10)
where $z_{\epsilon_{PE}/2}=\text{erf}^{-1}(1-\epsilon_{PE}/2)$ and $\text{erf}(x)$ is the error function.
Definition (9) and (10) guarantees that we never estimate a transmittance higher than its real value or a noise variance lower than its real value, except with probability $\epsilon_{PE}/2$. Thus, it holds both composability and operational interpretation.
In this context, the covariance matrix assuming the probability of failure of MLE is given by
$\Gamma_{\epsilon_{PE}}=\begin{pmatrix}(V_{A}+1)\mathbb{I}_{2}&t_{min}Z\sigma_{z}\ \\ t_{min}Z\sigma_{z}&(t_{min}^{2}V_{A}+\sigma^{2}_{max})\mathbb{I}_{2}\\ \end{pmatrix}\,$, (11)
which means that there exists a confidence set $\mathcal{C}_{\epsilon_{PE}}$ such that the covariance matrix $\Gamma_{\epsilon_{PE}}$ lies within $\mathcal{C}_{\epsilon_{PE}}$ with probability at least $1-\epsilon_{PE}/2$. Thus, the secret key rate that accounts for the probability of failure in parameter estimation can be computed using PEP 1.
Parameter estimation protocol 1
- Parameter estimation via maximum likelihood estimation in the finite-size scenario
  1. 1.
    Since Alice only has access to his measured signals $y$, Bob needs to broadcast $m$ signals over an authenticated channel so that Bob can estimate $t$ and $\sigma^{2}$.
  2. 2.
    Alice uses estimator from Eq. (8) to estimate $t$ and $\sigma^{2}$, using the $m$ correlated data.
  3. 3.
    Alice uses the statistical analysis from Eq. (9) to compute $t_{min}$ and from Eq. (10) to compute $\sigma^{2}_{max}$.
  4. 4.
    Alice uses these results to write the covariance matrix from Eq. (11) and, finally, compute $\chi_{\epsilon_{PE}}(y:E)$.
Note that step 3 becomes redundant in the asymptotic limit.
IV Worst-case confidence intervals for neural networks
The computational modeling of systems with output $Y$ is described by a function $f(\mathrm{X},\theta^{*})$, where $\theta^{*}$ denotes the parameters of the model. The output is assumed to be affected by an additive error term $\varepsilon$, which is independently and identically distributed according to a normal distribution $\varepsilon\sim\mathcal{N}(0,\sigma^{2}_{\varepsilon})$. For each observation $i=1,2,\ldots,N$, the model is represented as
$\mathrm{Y}_{i}=f(\mathrm{X}_{i},\theta^{*})+\varepsilon_{i}\,$, (12)
where $\mathrm{X}_{i}$ denotes the input corresponding to the $i$-th observation Seber and Wild (2005). This modeling framework is well established in computational learning theory, as it enables statistical generalizations across a broad class of inference tasks Barron (1994); Papadopoulos et al. (2001). The central challenge, therefore, lies in demonstrating that neural networks can be effectively described within this framework under appropriate assumptions, thereby ensuring reliable and secure parameter estimation. In this work, we adopt such a perspective, drawing inspiration from the delta method outlined in refs. Chryssolouris et al. (1996); Hwang and Ding (1997).
In general, neural networks for prediction give an output
[…]
which can approximate from Eq. (12) by minimizing the error function
$S(\theta)=\sum_{i=-1}^{N}[\hat{\mathrm{Y}}_{i}-f(\mathrm{X}_{i};\theta^{*})]^{2}\,.$ (14)
This procedure is expected to bring $\hat{\theta}$ closer to $\theta^{*}$. In principle, minimizing the loss function corresponds to finding the optimal set of network parameters that best approximates the underlying functional relationship between the inputs and the outputs Li et al. (2018). Given a sufficiently expressive architecture and representative training data, the neural network learns a map that minimizes the discrepancy between the predicted and true values of the target variable.
[…]
$f(\mathrm{X}_{i};\hat{\theta})\approx f(\mathrm{X}_{i},\theta^{*})+\mathbf{f}_{0}^{T}\cdot(\hat{\theta}-\theta^{*})\,$, (15)
where
$\mathbf{f}_{0}^{T}=\left(\frac{\partial f(\mathrm{X}_{i},\theta^{*})}{\partial\theta_{1}^{*}},\frac{\partial f(\mathrm{X}_{i},\theta^{*})}{\partial\theta_{2}^{*}},\ldots,\frac{\partial f(\mathrm{X}_{i},\theta^{*})}{\partial\theta_{p}^{*}}\right)\,$ (16)
with the subscript “0” indicating the set of points that are not used in the least-squares estimation of $\theta^{*}$. In this sense, the difference between the real and predicted value is written as
$\mathrm{Y}_{0}-\hat{\mathrm{Y}}_{0}\approx\varepsilon_{0}-\mathbf{f}_{0}^{T}\cdot(\hat{\theta}-\theta^{*})$ (17)
The first term corresponds to the intrinsic measurement noise, while the second term captures the uncertainty in the model prediction due to the estimation error in the parameters. The minimization process justifies modeling the noise $\varepsilon_{0}$ as a zero-mean Gaussian random variable with variance $\sigma^{2}_{\varepsilon}$, such that the parameter estimation error $\hat{\theta}-\theta^{*}$ can be approximated as following a multivariate normal distribution
$\hat{\theta}-\theta^{*}\sim\mathcal{N}_{p}\left(0,\sigma^{2}_{\varepsilon}\left[\mathbf{F}^{T}(\hat{\theta})\mathbf{F}(\hat{\theta})\right]^{-1}\right),$ (18)
[…]
$\mathbf{F}(\hat{\theta})=\frac{\partial f(\boldsymbol{\mathrm{X}},\hat{\theta})}{\partial\hat{\theta}}=\begin{bmatrix}\frac{\partial f_{1}(\mathrm{X}_{1},\hat{\theta})}{\partial\hat{\theta}_{1}}&\frac{\partial f_{1}(\mathrm{X}_{1},\hat{\theta})}{\partial\hat{\theta}_{2}}&\cdots&\frac{\partial f_{1}(\mathrm{X}_{1},\hat{\theta})}{\partial\hat{\theta}_{p}}\\ \frac{\partial f_{2}(\mathrm{X}_{2},\hat{\theta})}{\partial\hat{\theta}_{1}}&\frac{\partial f_{2}(\mathrm{X}_{2},\hat{\theta})}{\partial\hat{\theta}_{2}}&\cdots&\frac{\partial f_{2}(\mathrm{X}_{2},\hat{\theta})}{\partial\hat{\theta}_{p}}\\ \vdots&\vdots&\ddots&\vdots\\ \frac{\partial f_{n}(\mathrm{X}_{n},\hat{\theta})}{\partial\hat{\theta}_{1}}&\frac{\partial f_{n}(\mathrm{X}_{n},\hat{\theta})}{\partial\hat{\theta}_{2}}&\cdots&\frac{\partial f_{n}(\mathrm{X}_{n},\hat{\theta})}{\partial\hat{\theta}_{p}}\end{bmatrix}$ (19)
This formulation reflects how parameter uncertainty contributes to the overall variance of the model’s prediction, resulting in
$\mathrm{var}[\mathrm{Y}_{0}-\hat{\mathrm{Y}}_{0}]\approx\sigma^{2}_{\varepsilon}+\sigma^{2}_{\varepsilon}\mathbf{f}_{0}^{T}(\mathbf{F}^{T}\mathbf{F})^{-1}\mathbf{f}_{0}\,.$ (20)
[…]
$s^{2}=\frac{\left\|\mathbf{\mathrm{Y}}-\mathbf{f}(\boldsymbol{\mathrm{X}},\hat{\theta})\right\|^{2}}{n-p}\,$, (21)
which is constant for a given $\mathrm{Y}\in\mathcal{S}$, where $S$ delimits the set of possible values for the excess noise estimation. This variance estimator is then used in the Student’s $t$-distribution to quantify the uncertainty associated with the prediction $\hat{\mathrm{Y}}_{0}$. Thus, one can construct a confidence interval for the predicted value $\hat{\mathrm{Y}}_{0}$, which incorporates both the estimated variance and the sensitivity of the prediction to the parameters through the Jacobian vector $\mathbf{f}_{0}$:
$t_{n-p}\sim\frac{\mathrm{Y}_{0}-\hat{\mathrm{Y}}_{0}}{\sqrt{\mathrm{var}[\mathrm{Y}_{0}-\hat{\mathrm{Y}}_{0}]}}\approx\frac{\mathrm{Y}_{0}-\hat{\mathrm{Y}}_{0}}{\sqrt{s^{2}+s^{2}\mathbf{f}_{0}^{T}(\mathbf{F}^{T}\mathbf{F})^{-1}\mathbf{f}_{0}}}$ (22)
$\approx\frac{\mathrm{Y}_{0}-\hat{\mathrm{Y}}_{0}}{s\left(1+\mathbf{f}_{0}^{T}(\mathbf{F}^{T}\mathbf{F})^{-1}\mathbf{f}_{0}\right)^{1/2}}$
and finally
$\mathrm{Y}_{0}-\hat{\mathrm{Y}}_{0}\pm t_{n-p}^{\alpha/2}s\left(1+\mathbf{f}_{0}^{T}(\mathbf{F}^{T}\mathbf{F})^{-1}\mathbf{f}_{0}\right)^{1/2}\,.$ (23)
From a computational picture, the trained neural network is defined as a family of probability distributions on a sample space of excess noise $S$, indexed by a parameter vector $\hat{\theta}\in\Theta$. Thus, the trained neural network acts as a statistical estimator $\hat{\theta}:S\to\Theta$, approximating the mapping from data to channel parameters. The quantity in Eq. (24) defines a confidence interval with probability $1-\epsilon/2$, which is operationally equivalent to the parameter estimation method based on MLE in ref. Leverrier et al. (2010):
$\sigma^{2}_{max-NN}\approx\hat{\sigma}^{2}_{NN}+t_{n-p}^{\epsilon_{PE}/2}s(1+\mathbf{f}_{0}^{T}\left[\mathbf{F}^{T}(\hat{\theta})\mathbf{F}(\hat{\theta})\right]^{-1}\mathbf{f}_{0})\,.$ (24)
Since neural networks can be computationally expensive, the efforts invested in them must be used on processes that have significant impacts on the key rate. As discussed, this is the case for excess noise Huang et al. (2016); Laudenbach et al. (2018); Hajomer et al. (2024). In this case, the neural network to estimate the variance in the worst-case is then defined, and finally, we can estimate the excess noise via PEP 2, which is operationally equivalent to PEP 1.
Parameter estimation protocol 2
- Parameter estimation via neural network in the finite-size scenario
  1. 1.
    Since Alice only has access to his measured signals $y$, Bob needs to broadcast $m$ signals over an authenticated channel so that Bob can estimate $t$ and $\sigma^{2}$.
  2. 2.
    Alice uses estimator from Eq. (8) to compute $\hat{t}$ and a trained neural network to compute $\hat{\sigma}^{2}$.
  3. 3.
    Alice uses the statistical analysis from Eq. (9) to compute $t_{min}$ and from Eq. (24) to compute $\sigma^{2}_{max}$.
  4. 4.
    Alice uses these results to write the covariance matrix from Eq. (11) and, finally, compute $\chi_{\epsilon_{\mathrm{PE}}}(y:E)$.
[…]
V Neural network model
In this section, we present the neural network architecture developed to implement the framework introduced previously. We detail the data input structure, the network’s architectural design, and the training strategy employed. A discussion regarding the computational cost associated with the proposed neural network is provided in Appendix A.
V.1 Neural network inputs
To estimate the noise variance parameter in a CV-QKD system, we designed a fully-connected feedforward neural network tailored to extract nonlinear correlations from statistical quadrature measurements of the channel. All input data are expressed in shot-noise units (SNU), and the network is trained to estimate the product $\hat{t}^{2}\hat{\xi}$, since the parameter $\mu$ is constant for diferent frames.
The input vector to the network is given by $\mathrm{X}_{i}\in\mathbb{R}^{6}$, composed of sufficient statistics computed from a sample of correlated variables $\{x_{i},y_{i}\}_{m}$:
$\mathrm{X}_{i}=\left\{\hat{t}_{\text{MLE}},\ \langle x\rangle,\ \langle y\rangle,\ \text{Var}(x),\ a^{2}(\text{Var}(y^{\prime})-1),\ \text{Cov}(x,y^{\prime})\right\},$ (25)
where $\hat{t}_{\text{MLE}}$ is given by Eq. (8), and $y^{\prime}$ is a preprocessed version of Bob’s variable $y$, defined as:
$y^{\prime}_{i}=y_{i}-\hat{t}_{\text{MLE}}x_{i}+\frac{\hat{t}_{\text{MLE}}}{a}x_{i},$ (26)
with $a>1$ representing an artificial amplification factor. This preprocessing step is designed to enhance the contribution of excess noise in the signal, making it more detectable by the neural network.
Under this transformation, the variance of $y^{\prime}$ becomes:
$\text{Var}(y^{\prime})=(\hat{t}-\hat{t}_{\text{MLE}})^{2}V_{A}+\frac{\hat{t}_{\text{MLE}}^{2}}{a^{2}}V_{A}+\hat{t}^{2}\xi+1,$ (27)
and the rescaled quantity $a^{2}(\text{Var}(y^{\prime})-1)$ used as an input feature isolates the amplified noise components:
$a^{2}(\text{Var}(y^{\prime})-1)\approx\hat{t}_{\text{MLE}}^{2}V_{A}+a^{2}\hat{t}^{2}\xi,$ (28)
since the discrepancy $(\hat{t}-\hat{t}_{\text{MLE}})^{2}$ vanishes in the large-sample limit due to the consistency of the MLE.
The covariance term $\text{Cov}(x,y^{\prime})$ is computed from the sample using the standard Pearson correlation estimator. The output of the network is then post-processed by dividing by $a^{2}$, recovering an accurate estimate of the original parameter $\hat{t}^{2}\hat{\xi}$ from the amplified noise features.
V.2 Network Architecture
The architecture of the network is illustrated in Fig.  1. The main goal here was to test the framework using a simple architecture, which comprises:
  • An input layer with six entry points, each corresponding to one of the features in $\mathrm{X}_{i}$.
  • A first hidden layer with 32 neurons using ReLU (Rectified Linear Unit) activation.
  • A second hidden layer with 64 neurons, also using ReLU activation.
  • A third hidden layer with 32 neurons without an explicit activation function prior to the final output transformation.
  • An output layer consisting of a single neuron, whose output is passed through a shifted Softplus activation function, defined as:
    $\hat{\mathrm{Y}}=\log(1+e^{z+b}),$ (29)
    where $z$ is the output of the final hidden layer and $b\in\mathbb{R}$ is a learnable bias parameter initialized with a small positive value ($b=0.1$) to encourage strictly positive predictions.
Figure 1: Neural network architecture for estimating the scaled excess noise $t^{2}\xi$.
[…]
V.3 Training Strategy
The network is trained using supervised learning, where the target values are derived from synthetic data generated via an additive white Gaussian noise (AWGN) model (see Eq. (1)). The predicted output $\hat{\mathrm{Y}}$ corresponds to the scaled noise term $a^{2}\hat{t}^{2}\hat{\xi}$, allowing a later inversion to recover the physical excess noise. The loss function is computed using the mean square error to ensure the minimization of Eq. (14). The model is optimized using the Adamw optimizer Loshchilov and Hutter (2019) via the Optax library, leveraging the JAX and Flax frameworks for high-performance computation.
VI Numerical investigations
We investigate a finite-size security analysis employing neural networks within a specific architecture that leverages the signals $\{y_{i}\}_{m}$ and $\{x_{i}\}_{m}$ required for parameter estimation. A total of $10^{5}$ transmissivity values $t$ are sampled, each associated with a corresponding noise variance $\sigma^{2}=1+t^{2}\xi$. For every sampled pair, we generate computationally $N$ signal using the discussed protocol. Channel parameters are then estimated using $m\equiv N/2$ signals in the PEP 1 and PEP 2, which implement MLE and a neural network-based approach, respectively. This results in $10^{5}$ estimates for both $\sigma^{2}_{\text{max-MLE}}$ and $\sigma^{2}_{\text{max-NN}}$.
As an initial benchmark, we compare the precision of the estimators by analyzing the standard deviation between the predicted and true values of $\sigma^{2}$. As shown in Fig. 2, the neural network consistently yields lower deviations than the conventional MLE, reflecting the effectiveness of the error minimization performed during training. This increased accuracy stems from the fact that the network parameters are optimized to reduce a cost function, such as the mean squared error, which directly penalizes large prediction errors. Furthermore, one can verify that $\sigma^{2}_{\text{max}}\rightarrow\sigma^{2}$ as the sample size $m$ increases, as expected from asymptotic consistency of both estimators (see Eq. (10) and Eq. (24)).
Figure 2: Standard deviation between the estimated channel parameters $\sigma^{2}_{\text{max-NN}}$ and $\sigma^{2}_{\text{max-MLE}}$ and the real values $\sigma^{2}$, using $m=10^{4}$, $10^{5}$, $10^{6}$, $10^{7}$ and $10^{8}$ signals. The curves show that the average distance between the neural network estimation and the real values is smaller, which implies more precise estimations if compared to standard MLE method.
However, the main challenge is not simply to show that neural networks can be more precise (a result already discussed in literature Liu et al. (2018); Chin et al. (2021); Luo et al. (2022)), but demonstrate that they can also be $\epsilon_{PE}\text{-secure}$ for parameter estimation in CV-QKD. An estimate is $\epsilon_{PE}\text{-secure}$ if, and only if, all the points estimated are inside the confidence intervals with probability at least $1-\epsilon_{PE}/2$, i.e., one can never estimate $\sigma^{2}_{\text{max}}$ below its real value considering this probability. Figure 3 illustrates this behavior by depicting the average trend line obtained from the closest predicted points to the ideal reference, computed across all samples. This visualization emphasizes the proximity between predicted and true values. Notably, in all cases, the estimates remained above the ideal curve, ensuring that no underestimation occurred throughout the tested configurations.
Figure 3: Comparison between the estimated and real values of $\sigma^{2}$. Dot-dashed line, dashed line and dotted line corresponds, respectively, to $m=10^{6}$, $10^{7}$ and $10^{8}$ signals. In all cases, the estimated values was never inferior to the real values.
To analyze the impact of improved parameter estimation on the secret-key rate, we set $p_{ec}=0.9$ for the probability of successful error correction and adopt $\epsilon_{PE}=\epsilon_{\text{cor}}=\bar{\epsilon}=\epsilon_{PA}=10^{-10}$ for the security parameters, yielding an overall composable security level of $\epsilon\approx 3.9\cdot 10^{-10}$ against collective Gaussian attacks, as described in Eq. (6). The parameters summarized in Tab.  LABEL:tab:parameters are selected to reflect realistic conditions, based on experimental implementations reported in the literature Zhang et al. (2024); Lu et al. (2019); Milicevic et al. (2018); Chin et al. (2021); Jain et al. (2022), aiming to ensure practical feasibility.
Table 1: Protocol parameters used in this work. The parameter $d$ is chosen based on ref. Jain et al. (2022). Detector efficiency values reflect specifications of commercially available detectors Zhang et al. (2024); Lu et al. (2019). The reconciliation efficiency and probability of success of error correction is set according to recent experimental implementations Milicevic et al. (2018); Chin et al. (2021); Jain et al. (2022).
Protocol parameter Symbol Value
discretization $d$ 6
Quantum duty $\mu$ 1 (hom.)
Detector efficiency $\eta_{\mathrm{eff}}$ 0.8
Excess noise $\xi$ $0.01$ SNU
Variance $V_{A}$ 55 SNU
Left signals fraction $n/N$ $0.5$
Reconciliation efficiency $\beta$ $0.95$
Probability of success $p_{ec}$ $0.9$
of error correction
The precision of parameter estimation plays a critical role in the secret-key rate — since we must overestimate the excess noise with high confidence, more accurate estimators yield smaller values of $\xi$, ensuring that $\Gamma_{\epsilon_{PE}}\in\mathcal{C}_{\epsilon_{PE}}$. The results in Fig. 4 illustrate this behavior: the estimated secret-key rate consistently remains below the ideal rate, as expected. This outcome confirms the operational security of the parameter estimation procedure described in PEP 2, with the advantage of achieving higher rates.
Figure 4: Secret-key rate using the discussed protocol with parameters described in Tab. LABEL:tab:parameters. Dot-dashed line, dashed line and dotted line corresponds, respectively, to $N=2\cdot 10^{6}$, $2\cdot 10^{7}$ and $2\cdot 10^{8}$ signals. The use of neural networks allowed a gain of $6.1$ km, $1.3$ km and $3.0$ km, respectively. In all cases, the estimated values was never superior to the real values
Although the computational cost during training is considerable, inference is highly efficient and readily implementable in practical scenarios. This balance is particularly critical in the finite-size regime with limited signals, where even small improvements in parameter estimation can significantly extend the achievable communication distance between Alice and Bob, enabling secure key distribution in conditions where traditional estimators fall short.
VII Conclusion
Neural networks have been increasingly adopted as estimation tools in quantum information. In the context of QKD, the trade-off between computational cost and estimation accuracy becomes particularly relevant. While neural networks typically demand greater computational resources compared to conventional methods, they can offer improved precision. This is significant in the finite-size regime, where even modest gains in parameter estimation may translate into a positive key rate in otherwise insecure regimes.
In conclusion, this article gives a finite-size analysis for secure CV-QKD using networks for excess noise estimation. While the neural network employed in our simulations demonstrates improved estimation accuracy compared to the conventional MLE, we emphasize that the primary contribution of this work lies not in the superiority of a specific architecture, but in demonstrating that neural network-based estimators can be incorporated into parameter estimation routines for CV-QKD without compromising composable security. Although more robust or efficient architectures may be developed, our findings indicate that such data-driven approaches are compatible with finite-size security proofs. This insight enables the use of flexible and potentially adaptive estimation strategies in practical QKD systems, paving the way for further integration of machine learning techniques into secure quantum communication protocols.
Acknowledgements.
This work has been fully funded by the project Computational Architecture for Flexible QKD System Post-processing Platform supported by QuIIN - Quantum Industrial Innovation, EMBRAPII CIMATEC Competence Center in Quantum Technologies, with financial resources from the PPI IoT/Manufatura 4.0 of the MCTI grant number 053/2023, signed with EMBRAPII.
Appendix A Computational cost of the neural network
The computational complexity of the neural architecture is primarily determined by the number of trainable parameters and the per-sample inference cost. The model is implemented in Flax and trained using the JAX framework, leveraging XLA compilation and hardware acceleration for efficient execution. Let $n$ be the number of samples per iteration and $d$ the input dimension, with $d=6$ corresponding to the MLE estimate, mean values, variances, and covariances extracted from the AWGN model. The network processes input tensors of shape $(n,d)$.
Assuming a fully connected feedforward neural network with $L$ layers and $h$ hidden units per layer, the time complexity of a forward or backward pass is approximated by:
$\mathcal{O}(dh+(L-1)h^{2})\approx\mathcal{O}(Ldh+Lh^{2})\,.$ (30)
The term $\mathcal{O}(dh)$ corresponds to the affine transformation from the input to the first hidden layer, while $\mathcal{O}((L-1)h^{2})$ results from the matrix multiplications between subsequent hidden layers Goodfellow et al. (2016).
Training and inference benefit from JAX primitives such as jit and vmap, enabling automatic parallelization and just-in-time compilation. Pseudo-random number generation for AWGN simulations is handled deterministically via key splitting to ensure reproducibility.